Prev Next

SW&IT Meeting Minutes 2021-02-08

Attendees: David, Igor, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20210125
LR20210125 - Bec, Matthieu 80.0
LR20210125 - Edmeades, David 80.0
LR20210125 - Sola, Igor 24.0
LR20210125 - Grenz, Paul 0.0
LR20210125 - Hooper, Stephen 0.0
LR20210125 - Zhang, Yang 80.0
LR20210125 - Zhang, Xianyu 77.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • ATT S/W review slides
    • AGW OAC encoder update meeting
    • Web student posting
  • This week
    • ATT S/W slides complete, AORR review
    • SHARK NIR RFAs
    • HBS plc upgrade discussion
    • AGW hands-on with JP
    • DMS react.js
    • Web student review candidates

DE

  • Archive stuff
    • Greenlight from Cristina for shutting archive.mountain down
    • Nothing so far regarding directory stuff, I'll keep pushing
  • Cobbler web redirect stuff: SH
  • New storage doc
  • Office visit Monday
    • Any tasks let me know
  • Another reminder: Saturday power transfer; unknown if our equipment will stay up long enough to make it to generator power

XZ

  • Last week
    • ATT AdSec and SW feedback from reviewer
    • FSM Fault recovery doucment
    • Analyze one SX AdSec DSP watdog exipre event
    • AO support
  • This week
    • Work on SX AdSec issues
    • Track SOUL issue
    • Friday off
    • AO support

YZ

  • Previous:
    • Some TCS work (IIF patch release, still revising the TCS docs).
    • Spent most time on TMS (meeting, software revision done, software spec released).
  • Next:
    • TCS doc revision to continue.
    • Possible TMS software test sometime (not sure when yet).
    • Will mainly focus on OSCO software.

SH

  • Steward meeting
  • UITS meeting
  • core-b work (LeRoy handled it)
  • archive
  • e-mail groups

next:

  • Steward meeting -- power down Saturday/Monday
  • DIMM
  • robs2
  • UITS WNR stuff

PSG

  • Last week
    • Joined new UMAC project
    • Updated OSS code, merged swing arm.
  • This week
    • Continuingto update OSS code.

IS

Last week:
  • Took most of the week off sick

This week:
  • LUCI
    • IT #7493 Aladin slow subtraction
    • IT #8231- Timeout on GEIRS during readout
    • IT #8026- Possibly implementing Update OT with new AO patrol field parameters.
    • create a luci rpm via a docker container (sorry! I keep forgetting)
Topic revision: r9 - 08 Feb 2021, IgorSola
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback